OpenStudy (anonymous):

multiple choice question! please help! Suppose y varies directly with x, and y = 14 when x = –4. What is the value of y when x = –6? –21 10 18 21

jimthompson5910 (jim_thompson5910):

y varies directly with x means that y = kx

jimthompson5910 (jim_thompson5910):

when x = -4, y is y = 14 so y = kx turns into 14 = k*(-4)

jimthompson5910 (jim_thompson5910):

solve for k and tell me what you get

OpenStudy (anonymous):

-3.5, But I think Im wrong..

jimthompson5910 (jim_thompson5910):

you're correct

jimthompson5910 (jim_thompson5910):

so y = kx turns into y = -3.5x

jimthompson5910 (jim_thompson5910):

now they want to know the value of y when x = -6

jimthompson5910 (jim_thompson5910):

so y = -3.5x y = -3.5(-6) y = ???

OpenStudy (anonymous):

21.

OpenStudy (anonymous):

thank yoouuu!:))

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (anonymous):

can you walk me through this one too? Suppose y varies directly with x, and y = 25 when x = 140. What is the value of x when y = 36?

jimthompson5910 (jim_thompson5910):

y varies directly with x means what

OpenStudy (anonymous):

y=kx

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we're given x = 140 and y = 25 so y = kx becomes 25 = k*140

jimthompson5910 (jim_thompson5910):

so k = ____

OpenStudy (anonymous):

.18

jimthompson5910 (jim_thompson5910):

I would use more decimal places for k. I'm going to use k = 0.17857 so y = kx becomes y = 0.17857x Now plug in y = 36 and solve for x y = 0.17857x 36 = 0.17857x 36/ 0.17857 = x 201.601613 = x x = 201.601613 now round to the nearest tenth to get 201.6 this is the same as 201 and 3/5 or the improper fraction 1008/5

jimthompson5910 (jim_thompson5910):

optionally... y = kx 25 = k*140 25/140 = k k = 25/140 k = 5/28 -------------------- y = kx y = (5/28)x 36 = (5/28)*x 36*28 = 5x 1008 = 5x 5x = 1008 x = 1008/5 So x = 1008/5 which is x = 201.6 in decimal form